World Enterprises is determined to report earnings per share of $2.05. It therefore acquires the Wheelrim and Axle Company. You are given the following facts:

World
Enterprises
Wheelrim
and Axle
Merged Firm
Earnings per share $ 1.50 $ 2.00 $2.05
Price per share $ 30.00 $ 20.00 ?
Price–earnings ratio 20 10 ?
Number of shares 110,000 200,000 ?
Total earnings $ 165,000 $ 400,000 ?
Total market value $ 3,300,000 $ 4,000,000 ?


There are no gains from merging. In exchange for Wheelrim and Axle shares, World Enterprises issues just enough of its own shares to ensure its $2.05 earnings per share objective.

Required:

  1. Complete the below table for the merged firm.
  2. How many shares of World Enterprises are exchanged for each share of Wheelrim and Axle?
  3. What is the cost of the merger to World Enterprises?
  4. What is the change in the total market value of the World Enterprises shares that were outstanding before the merger?
Answer #1
Q1. Complete the above table.
Ans.
World Enterprises Wheelrim and Axle Merged Firm
Earnings per share $1.50 $2 $2.05
Price per share $30 $20 $26.487   (refer note 4)
Price–earnings ratio 20 10 12.92     (refer note 5)
Number of shares 110,000 200,000 275,610 (refer note 2)
Total earnings $165,000 $400,000 $565,000 (refer note 1)
Total market value $3,300,000 $4,000,000 $7,300,000 (refer note 3)
Note 1 - As there are no gains from merging the total earnings of the merged firm would be $165,000 + $400,000 = $565,000
Note 2 - Total number of shares of the merged firm = Total Earnings / Earnings per share
= 565,000/2.05
= 275609.756
= 275610
Note 3 - Total market value of the merger firm = $3,300,000 + $4,000,000 = $7,300,000
Note 4 - Price per share of the merged firm = Total Value of the merged firm / Number of shares of the merged firm
= $7300000/275610
= 26.487
Note 5 - Price Earning Ratio = Price per share/ Earnings per share
= 26.487/2.05
= 12.92
Q2.
How many shares of World Enterprises are exchanged for each share of Wheelrim and Axle?
Ans.
Total number of shares of merged firm = 275,610
Shares held by shareholders of World enterprises = 110,000
Shares issued to shareholders of Wheelrim and Axle= 275,610-110,000
=          165,610
Exchange ratio = 165610/ 200000
= 82.8%
Q3. What is the cost of the merger to World Enterprises?
Ans.
Cost of merger = No of shares to be issued to vendor company/ Total number of shares of merged company * Value of merged firm - Value of Vendor firm
= 165610/275610*7,300,000- 4,000,000
= $386,463
Q4. What is the change in the total market value of the World Enterprises shares that were outstanding before the merger?
Ans.
Change in Total market value of the shareholders of World Enterprises shares that were outstanding before the merger = 110,000 * (30-26.487)
=              386,430
Value of the shares that were outstanding before merger have been reduced by $386,430 after merger